OpenStudy (anonymous):

How would I write a function with a phase shift of 90degrees? Or more over, how do I calculate the phase shift?

OpenStudy (dan815):

t=sin(x-p) p= phase shift p=pi/2 rad or 90 degrees

OpenStudy (anonymous):

Okay thanks! So is pi 180 degrees?